OpenStudy (anonymous):

help!??! A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 27π in3 65π in3 105π in3

OpenStudy (anonymous):

@rajat97 help plz

OpenStudy (rajat97):

hello i'm here

OpenStudy (rajat97):

and first of all i wanna wish you a merry merry christmas to you

OpenStudy (anonymous):

thx and i wish u a Merry Christmas too

OpenStudy (rajat97):

:) okay first of all you need to find the volume of the whole vase next, find the volume of one marble and then multiply it by 9 next, what you need is the volume of the water that is poured in the vase so the volume of the water will be equal to the remaining volume so, volume of liquid=volume of cylinder-TOTAL volume of marbles i'll post the working for you

OpenStudy (rajat97):

diameter(d) of cylinder=6inches so radius of cyl=r=3inches and the height of the cyl=h=12 inches (h=12 inches as the water is filled upto 12 inches and we'll take into consideration only the cylinder upto that height i hope you understand this) so the volume of the cyl=\[\Pi r^2 h\] =pi x 3 x 3 x 12 =108pi next, the volume of one sphere = 4/3 x pi x r^3 (here , d=2 so , r=1) =4/3 x pi x 1 x 1 x 1=4pi/3 so the volume of water will be 108pi-4pi/3 x 9 (9=total no. of marbles) =96pi that is first option

OpenStudy (anonymous):

thx a lot

OpenStudy (rajat97):

you're welcome and thanks for the medal

OpenStudy (anonymous):

u are wlc